A store owner plant to make 10 pounds of a candy mix worth 1.95/lb. How many pounds of gummy bears worth $2.40/lb and how many pounds of jelly beans worth $1.60/lb must be combined to make the candy mix?
Answer by ptaylor(2198)

Let x=amount of gummy bears needed
($understood)
Value of the gummy bears=2.40x
Then 10-x=amount of jelly beans needed
Value of the jelly beans=1.60(10-x)=16-1.60x
Now we know that the value of the gummy bears (2.40x) plus the value of the jelly beans(16-1.60x) has to equal the value of the final mixture (1.95*10). So, our equation to solve is:
2.40x+16-1.60x=1.95*10 subtract 16 from each side and simplify
2.40x+16-16-1.60x=19.5-16 collect like terms
0.80x=3.5 divide each side by 0.80
x=4.375 lbs-----amount of gummy bears needed
10-x=10-4.375=5.625 lbs-----amount of jelly beans needed
CK
2.40*4.375 +1.60*5.625=1.95*10
10.5+9=19.5
19.5=19.5

Hope this helps---ptaylor